Mujtaba covered 500m in 30 minutes find the speed in metre per second and km per hour

1
Expert's answer
2021-12-12T16:43:32-0500

By definition, the speed is:


"v = \dfrac{500m}{30min} = \dfrac{50}{3}m/min"

In meters per second:


"v = \dfrac{50}{3}\cdot \dfrac{1}{60}m/s \approx 0.28m/s"

In km/h:


"v = \dfrac{50}{3}\cdot 0.001\cdot 60 = 1km/h"